OpenStudy (anonymous):

What is the slope of the line that passes through the points (-2,5) and (1,4)?

OpenStudy (mathstudent55):

The line through points \((x_1, y_1)\) and \((x_2, y_2) \) has slope: \(m = \dfrac{y_2 - y_1}{x_2 - x_1} \)

OpenStudy (mathstudent55):

Use the formula with your points.

OpenStudy (anonymous):

I have no idea what you mean

OpenStudy (mathstudent55):

Ok. I'll explain. We call the coordinates of one points x1 and y1. In your case you have 2 points. Let's call point (-2, 5) our first point. Then x1 = -2 and y1 = 5

OpenStudy (mathstudent55):

The second point is (1, 4). Then x2 = 1 and y2 = 4

OpenStudy (mathstudent55):

Remember that each point has 2 coordinates, the x-coordinate and the y-coordinate. For the first point, (-2, 5), the x-coordinate is -2, and the y-coordinate is 5. For the second point, the x-coordinate is 1, and the y-coordinate is 4.

OpenStudy (mathstudent55):

The reason we call the coordinates x1, y1, x2, y2 is just to keep track of which points they belong to. x1 is the x-coordinate of the first point. x2 is the x-coordinate of the second point. y1 is the y-coordinate of the first point. y2 is the y-coordinate of the second point.

OpenStudy (mathstudent55):

y2 - y1 means subtract the y-coordinate of the first point from the y-coordinate of the second point. Since y2 = 4 and y1 = 5, y2 - y1 = 4 - 5 = -1

OpenStudy (anonymous):

is the answer -3?

OpenStudy (mathstudent55):

x2 - x1 means subtract the x-coordinate of the first point from the x-coordinate of the second point. Since x2 = 1 and x1 = -2, y2 - y1 = 1 - (-2) = 1 + 2 = 3

OpenStudy (anonymous):

or -2?

OpenStudy (mathstudent55):